OpenStudy (anonymous):

suppose lim[x:4,f(x)=4] and lim[x:4,g(x)= -2], find, lim[x:4,g(x)/f(x)-1]

OpenStudy (anonymous):

use quotient rule for taking derivative.

OpenStudy (anonymous):

This appears to be a problem of limits. The very question states that the limits f(x)=4 and g(x)=-2 as x approaches 4 means we can assume those limits exist. Because we know: lim f(x)/g(x) = lim f(x)/lim g(x) .... the question becomes a simple substitution problem. If lim g(x)=0, the problem wouldn't have a defined solution. however our answer will be: lim[x:4, f(x)/g(x) -1] =(4)/(-2) -1 =-2-1 =-3